OpenStudy (anonymous):

Alfred graphed two lines, y = x + 4 and y = x - 2. How would this be graphed?

OpenStudy (cwrw238):

you could just plot two points for each line first one let x = 0 then y = 0+ 4 = 4 point(0,4) let x = 4 then y = 4+ 4 = 8 point (4,8) draw a straight line between the two points do similar procedure on second line